OpenStudy (amistre64):

internet is slow for me today .... Given 10 cards; 3 are red and 7 are blue. What is the probability of randomly drawing 5 cards and 2 of them are red?

OpenStudy (anonymous):

exactly 2 of them are red?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

find\[\frac{10!}{5!(10-5)!}\] which is 252... out of 252 sets of 5 cards.... how many of them are red?

OpenStudy (anonymous):

oops I mean out of those 252 sets how many of them contain exactly 2 reds?

OpenStudy (amistre64):

good, we have a total number of outcomes to be 252; of those outcomes, what is the chance of drawing 2 red cards and 3 blue?

OpenStudy (anonymous):

lemme do a python simulation

OpenStudy (amistre64):

so far we have: \[\frac{?}{^{10}C_5}\]

OpenStudy (anonymous):

so.... we find \[\frac{252!}{2!(252-2)!}\]?

OpenStudy (amistre64):

not quite ...

OpenStudy (amistre64):

or maybe lol

OpenStudy (anonymous):

draw 5 cards and you want exactly 2 red ones right?

OpenStudy (anonymous):

252! is a big number :-P

OpenStudy (amistre64):

it might be useful to think of it as needing \(^3C_2\) for your successes

OpenStudy (amistre64):

yes; draw 5 and exactly 2 are red

OpenStudy (anonymous):

we can probably reason it out, and get a nicer answer than this compound fraction, but it would be \[\frac{\dbinom{3}{2}\dbinom{7}{3}}{\dbinom{10}{5}}\]

OpenStudy (amistre64):

yes, a hypergeometric pdf :)

OpenStudy (anonymous):

why wouldn;t it be\[\frac{\dbinom{252}{2}}{\dbinom{10}{5}}\]

OpenStudy (amistre64):

its a related cousin to the binomial pdf

OpenStudy (anonymous):

reasoning for this as you need 2 out of 3 red, 3 out of 7 blue, divided by number of ways to pick 5 from ten

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

thats cool

OpenStudy (anonymous):

how do we do a python simulation to find an answer like this?

OpenStudy (amistre64):

10c5 outcomes total the number of successes would be 3c2 and failures 7c3

OpenStudy (amistre64):

binomial pdf relates to probability WITH replacement the hyper pdf relates to probability WITHOUT replacement

OpenStudy (anonymous):

i am not fond of these compound fractions though. we could also reason as follows: first card red with probability \[\frac{3}{10}\] second card red give first card red is \[\frac{2}{9}\] and then 3rd 4th and 5th blue given first stuff is \[\frac{7}{8}\times \frac{6}{7}\times \frac{5}{6}\] multiply all this together, and then multiply by number of ways to choose 2 from 5 and you should get the same answer

OpenStudy (anonymous):

my calculator says 5/12

OpenStudy (anonymous):

in other words (and i will check this number in case i am full of beans) you should also get \[\dbinom{5}{2}\times \frac{3\times 2\times 7\times 6\times 5}{10\times 9\times 8\times 7\times 6}\]

OpenStudy (anonymous):

i got \[\frac{5}{12}\] for first number, let me try second one

OpenStudy (amistre64):

i get 1/24

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

1/24 multiplied by binom(5,2)

OpenStudy (zarkon):

5/12 is correct

OpenStudy (amistre64):

5/12 is the right one yes

OpenStudy (anonymous):

it is \[\frac{1}{24}\times \dbinom{5}{2}\] what adg said

OpenStudy (anonymous):

just as a matter of taste i prefer the "reason it out" method. of course the numbers are the same after you invert, cancel etc, but i am not a fan of compound fractions like \[\frac{\dbinom{3}{2}\dbinom{7}{3}}{\dbinom{10}{5}}\]

OpenStudy (anonymous):

and not to belabor the obvious, but after the onslaught of cancellation this is a fairly simple arithmetic problem

OpenStudy (amistre64):

:)